OpenStudy (anonymous):

If a circle has a circumference of 64 pie inches, calculate the area of this circle. Please show work.

OpenStudy (anonymous):

\[2\pi r=64\] r=\[64/2\pi\] area = \[\pi r ^{2}\] =\[1024/\pi\]

OpenStudy (anonymous):

@djm it is pi as in the greek letter, not pie as in the dessert. just thought i would mention it

OpenStudy (anonymous):

circumference is 2*pi*r, so 2*pi*r = 64 * pi --> r = 32 Area is pi*r^2 = pi*32^2 = 1024*pi

OpenStudy (anonymous):

Haha, sorry satellite, I'm stressing for my final tomorrow and I can't focus...

OpenStudy (anonymous):

sorry i did'nt see the pie (pi) out there

OpenStudy (anonymous):

LOL